OpenStudy (anonymous):

Evaluate the expression without using a calculator: (-8)^1/3

Directrix (directrix):

(-8)^1/3 =[ (-2)^3)] ^ (1/3) = (-2) ^ [(3)(1/3)] = (-2)^ (1) = -2

OpenStudy (anonymous):

Thankyou very much for explaining it to me:)

Directrix (directrix):

You are welcome.